myininaya (myininaya):

satellite's question: Let n be positive integer. If we divide n by 3, the remainder is 2. If we divide n by 5, the remainder is 1. What is the smallest possible value for n?

OpenStudy (anonymous):

i get 11

OpenStudy (anonymous):

11

myininaya (myininaya):

lol gj

OpenStudy (anonymous):

3|n-2 5|n-1 n-2=3*q n-1=5*r 3q+2=5r+1 3q+1=5r so the smallest integers so this is true are q=3 r=2 that is n=11

OpenStudy (anonymous):

yep pretty similar to what I did

myininaya (myininaya):

how did you do it neel?

OpenStudy (anonymous):

number theory is pretty fun

myininaya (myininaya):

yes i agree

OpenStudy (anonymous):

i get to take a grad course in it starting in a few weeks

myininaya (myininaya):

i wish i could had taken a advanced number theory course :(

myininaya (myininaya):

none of the professors specialized in it

OpenStudy (anonymous):

we just have one who does so he teaches all the number theory

OpenStudy (anonymous):

and we have a really good one in abstract who's doing galois theory in the fall

myininaya (myininaya):

i really don't think i have the personality for math anymore since i only really enjoy number theory i should probably look into computer science or cryptography

OpenStudy (anonymous):

have you taken abstract algebra?

myininaya (myininaya):

yes

OpenStudy (anonymous):

if so you could get into coding

OpenStudy (anonymous):

thats mathematical comp science

myininaya (myininaya):

right i want to work at the nsa i think that might be cool but the location sucks

OpenStudy (anonymous):

what degree do you have in math?

myininaya (myininaya):

masters

OpenStudy (anonymous):

if you want to try a fun integral I posted one that no one tried: http://openstudy.com/groups/mathematics#/groups/mathematics/updates/4de848014c0e8b0bd934bcd8

myininaya (myininaya):

ok i will look at it